A particle starts moving with acceleration 2 ms-2 . distance travelled by it in the fifth half-second is:
1)1.25m
2)2.25m
3)6.25m
4)30.25m

Grade:11

2 Answers

bharat bajaj IIT Delhi
9 years ago
Distance travelled in fifth half second = Distance travelled in 5/2 seconds - Distance travelled in 2 seconds...s = 1/2. 2 . (5/2)^2 = 1/2.2.(2)^2 = 25/4 - 4 = 6.25 - 4 = 2.25 m..Option 2) is correct...THanks
Bharat BajajIIT Delhi
askiitians faculty
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty